# Securing Federated Learning Against Novel and Classic Backdoor Threats During Foundation Model Integration

### v1

@@ 4.1 Experimental Setup
Table 1: Defenses against the novel backdoor attack in homo-FL. D1: CIFAR-10, D2: CIFAR-100
@@
Table 2: Defenses against the novel backdoor attack in hete-FL. D1: CIFAR-10, D2: CIFAR-100
@@
Table 3: Proposed defense against classic backdoor attacks on CIFAR-10 in cross-silo homo-FL.
4.2 Experimental results
@@
Against Classic Backdoor Attack. Table 3 demonstrates the effectiveness of our defense method against classic backdoor attacks, including BadNet, Blend, and SIG, in the homo-FL cross-silo setting on CIFAR-10. To ensure the attacks remain potent, we compromise two clients and adjust the attack hyperparameters, resulting in most ASRs above 85% for vanilla models. In most cases, our method reduces the ASR by more than 50%, while keeping the ACC drop below 10%. The method demonstrates robustness across various attack strategies and FL settings, highlighting its applicability in defending against a wide range of backdoor threats in FL systems.
Applied in non-attack setting. Since our defense method does not assume the presence of an attack, we also evaluate its performance in clean (attack-free) settings, as shown under “No Attack” in Table 3. The drop in ACC is within 2% in the homogeneous cross-silo FL setting, demonstrating that our method minimally impacts model performance in the absence of an attack.
5 Conclusion

### v2

@@ IV Experiment
TABLE II: Defenses against the novel backdoor attack in hete-FL. D1: CIFAR-10, D2: CIFAR-100
@@
IV-A Experimental Setup
TABLE III: Proposed defense against classic backdoor attacks on CIFAR-10 in cross-silo homo-FL.
@@ IV-A Experimental Setup
IV-B Experimental Results
Against Novel Backdoor Attack. The vanilla FM-FL system demonstrates a high susceptibility to the novel backdoor attack across all FL settings and datasets. As shown under “Vanilla” in Table I and Table II, the ASRs exceed 90% in all configurations, while the ACC on clean samples remains at reasonable levels, highlighting the stealthy nature and effectiveness of this attack.
Existing defense methods, including NormThr, DP, Krum, ClipCluster, SignGuard, RFOUT, and Pruning, exhibit limited effectiveness against the novel threat. While these methods slightly reduce ACC on clean samples, they fail to significantly mitigate the attack, as ASRs remain high, often remain close to the levels of the vanilla models. This verifies that existing approaches, which rely heavily on identifying anomalies in client updates, are ill-suited to counter novel backdoor attacks that stem from poisoned synthetic data on the server side.
@@ IV-B Experimental Results
Against Classic Backdoor Attack. Table III demonstrates the effectiveness of our defense method against classic backdoor attacks, including BadNet, Blend, and SIG, in the homo-FL cross-silo setting on CIFAR-10. To ensure the attacks remain potent, we compromise two clients and adjust the attack hyperparameters, resulting in most ASRs above 85% for vanilla models. In most cases, our method reduces the ASR by more than 50%, while keeping the ACC drop below 10%. For BadNet, Blend, and SIG, ASR decreases substantially in IID settings with minimal impact on ACC. In non-IID settings, ASR drops to approximately 35%, similarly maintaining minimal impact on ACC. These results demonstrate robustness across various attack strategies and FL settings, highlighting its applicability in defending against a wide range of backdoor threats in FL systems.
Applied in non-attack setting. Since our defense method does not rely on the assumption of an attack, we evaluate its performance in clean (attack-free) settings to ensure it does not degrade model accuracy unnecessarily. As shown under “No Attack” in Table III, the drop in ACC remains within 2% in the homogeneous cross-silo FL setting. This minimal impact demonstrates the effectiveness of our defense in maintaining model performance on clean data. Furthermore, the negligible drop in ACC highlights the method’s practicality for deployment in real-world FL systems, where attacks may not always occur but defenses must remain active to ensure security without compromising model functionality.
V Conclusion